OpenStudy (anonymous):

Is a triangle with sides of 6, 8 and 10, a right triangle?

OpenStudy (anonymous):

no because 6^2+8^2 doesn't equal 10^2

OpenStudy (anonymous):

yes, if u use the formula a^2+b^2 = C^2 it fits with in this case it being 36+64=100 which is correct making it a right triangle ( because t hat formula only works for right triangles).

OpenStudy (anonymous):

sorry added wrong 36 + 64 = 100

OpenStudy (anonymous):

so yes

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks!